OpenStudy (anonymous):

Given that f(x) = x2 + 2x + 3 and g(x) = quantity of x plus four, over three, solve for f(g(x)) when x = 2.

OpenStudy (anonymous):

Same as the last one, but now g(x)=\[\frac{ x+4 }{ 3 }\], so you have to put that into f(x), solve the expression, and replace x with 2

OpenStudy (anonymous):

Is it 2 then?

OpenStudy (anonymous):

So f(g(x)) = ((x+4)/3)^2+2((x+4)/3) + 3

OpenStudy (anonymous):

and when x = 2, that should give (6/3)^2+2(6)+3

OpenStudy (anonymous):

So unfortunately the answer isn't 2

OpenStudy (anonymous):

Well when I answer it like that I get 19 and thats not one of the possible answers..

OpenStudy (anonymous):

Can you show me your working out? Because if you continue from the equation \[(6/3)^2+2*6+3\] then your answer shouldn't be 19

OpenStudy (anonymous):

remember PEMDAS

OpenStudy (ranga):

Whenever f(g(x)) is asked for a specific value of x, say x = 2, the easier approach is to find g(2) first and then put that value for x in f(x).

OpenStudy (anonymous):

Well I did the (6/3) and got 2 then I raised that to the second power and got 4. Then I multiplied 2 by 6 got 12 and added that to the 4 so I got 16 and then finally I added 3... So thats how I got 19.

OpenStudy (anonymous):

Oh damn you are right, I miscalculated :P

OpenStudy (anonymous):

The correct equation should be \[(6/3)^2+2*2+3\]

OpenStudy (anonymous):

Which should give 11

OpenStudy (anonymous):

Yup thats what I got now. And this one is a possibility! YAY!

OpenStudy (anonymous):

Great :D

OpenStudy (ranga):

g(x) = (x+4)/3 g(2) = 6/3 = 2 f(x) = x^2 + 2x + 3 f(g(2)) = f(2) = 2^2 + 2(2) + 3 = 11

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Don't sweat it